The equations in an economy are given as: C=Yd, Investment=320, Tax=300, Government expenditure=300, Exports=300-0.05Y.

a)Find the equilibrium level of income

b)Find the net exports at equilibrium level of income


1
Expert's answer
2021-01-08T12:58:30-0500

a) The consumption function is "C=Yd"

"C=(Y-T)"

"C=(Y-300)"

The equilibrium condition is given as "Y=C+I+G+X-M"

Thus, "Y=(Y-300)-0.05Y-0"

"Y=Y--0.05Y"

"Y-0.8YY=940"

"Y=940\/0.25"

The equilibrium level of income is "3760"

b) Imports "M=0" Net Exports

"X-M=300-0.05(3760)-0=300-188=112"